OpenStudy (anonymous):

sOLVE (2x+5)^3-(2x+1)^3=16

OpenStudy (harsimran_hs4):

start with put 2x + 3 = t (say) so eq becomes (t+2)^3 - (t - 2)^3 = 16 now solve for t

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is this correct

OpenStudy (harsimran_hs4):

(a+b)^3 = a^3 + b^3 + 3ab(a+b)

OpenStudy (harsimran_hs4):

(a-b)^3 = a^3 - b^3 - 3ab(a - b)

OpenStudy (anonymous):

I am stuck here 12y^2=0

OpenStudy (harsimran_hs4):

y is t ?

OpenStudy (harsimran_hs4):

if yes then y = 0 is the solution to the above equation

OpenStudy (anonymous):

yup

OpenStudy (harsimran_hs4):

2x+3 = t = 0 x = -3/2 this should be the answer

OpenStudy (anonymous):

yes the answer is S.S = {-3/2} I checked the answers

OpenStudy (harsimran_hs4):

cool :)